# Beyond Formula Complexity: Effective Information Criterion Improves Performance and Interpretability for Symbolic Regression

### v1

@@ 4 Experiments
In this section, we demonstrate practical applications of the Effective Information Criterion (EIC) in symbolic regression, including: 1) identifying prevalent unreasonable structures in existing SR results, 2) enhancing the performance of various heuristic search-based methods, 3) improving the sample efficiency of various generative methods, and 4) evaluating formula interpretability in alignment with domain expert intuition.
4.1 EIC Evaluates Existing Symbolic Regression Methods
@@
Experimental setups. To assess the prevalence of unreasonable structures in existing SR outputs, we compare the EIC values of 17 symbolic regression methods on 133 SRbench white-box problems against the ground-truth formulas. The SRbench set includes 119 Feynman physics formulas and 14 Strogatz ODE formulas, with corresponding synthetic data. Samples are randomly split into 75% training data for search and 25% test data for evaluation. Each method is run 10 times per problem, and results with $R^{2}>0.8$ are retrained, excluding poorly fitting formulas. This filtering removes approximately 39% of the data points.
Figure 4: Comparison of complexity and EIC between ground-truth physical formulas and symbolic regression results.
Finding 1. Gap with real-world formulas in terms of EIC. Figure 4 shows the average EIC values and complexities (formula lengths) of the evaluated algorithms alongside the 133 ground-truth formulas. Although most methods can discover compact formulas comparable in complexity to the ground truth, their EIC values are consistently higher, revealing widespread unreasonable structures and highlighting the need to incorporate EIC into evaluation.
Finding 2. Unit constraints and hand-designed rules improve structural plausibility. Among all methods, DSR achieves the closest EIC to the ground-truth, likely due to its use of dimensional constraints, which prevent nonlinear unary operators such as expexp and sinsin from being applied to dimensional subformulas, thereby pruning the search space and avoiding overly complex structures. PySR and GPlearn follow it with suboptimal EIC values, possibly because they restrict deep nesting of nonlinear unary functions through hand-designed rules, thus reducing structurally complex results.
Finding 3. Divide-and-conquer improves structural pausibility. AIFeynman2 and SR4MDL achieve performance similar to PySR and GPlearn, likely because they use a divide-and-conquer strategy that breaks the target formula into simpler subformulas, reducing abnormal structures. Notably, although AIFeynman2 produces formulas of high average complexity, their EIC values remain relatively low. This is likely due to its design, which uses a series of pre-defined symmetry rules to recursively rewrite formulas into smaller components. Since each divide-and-conquer step stems from simple symmetry rules, it avoids the unreasonable structures with high EIC values.
Finding 4. Generative methods tend to produce structurally unreasonable formulas. Compared to search-based methods, recent generative methods such as NeurSR, E2ESR, and SNIP yield formulas with much higher EIC values. This likely stems from training on randomly generated formulas that ignore structural plausibility, making them reproduce these abnormal substructures.
These findings highlight the effective practices in existing symbolic regression methods, while also revealing a gap between the structural plausibility of generated formulas and that of truly interpretable symbolic models, motivating our exploration of EIC to improve existing SR algorithms in the following experiments.
@@ 4.2 Incorporating EIC into Heuristic Search Methods
Experimental setups We evaluated GP and MCTS with and without EIC integrated on SRBench, which is a standard benchmark for SR, including 133 white-box problems with known ground-truth formulas and 122 black-box problems without explicit underlying dynamics. For each problem, we ran 10 independent trials with a four-hour time limit and compared results against the 17 baseline methods in Section 4.1. Data were split into 75% training and 25% test sets. For white-box problems, we tested both noise-free data and noisy data with three levels $\eta=0.001,0.01,0.1$, where Gaussian noise with standard deviation $\sigma=\eta\times\mathrm{Std}[y]$ was added to the outputs, allowing us to assess robustness under varying noise. Performances of algorithms are evaluated from two complementary perspectives, including 1) the trade-off between accuracy ($R^{2}$) and complexity (formula length) that reflects the ability of the algorithm to learn compact yet accurate models, and 2) the average EIC values of discovered formulas that reflect their structural plausibility.
White-box experimental results. Figure 5 summarizes the white-box results. As shown in the left panel, integrating EIC into MCTS and GP improved both accuracy and compactness, moving them from the second to the first tier of the Pareto front. This is because EIC, as a regularizer, can prevent overfitting and enhance generalization accuracy to test set data. It also reduces formula length by penalizing unnecessary complexity. In contrast, while methods such as DSR and RSRM produce compact formulas with relatively low EIC values (Figure 4), their predictive accuracy is much lower. This suggests that although unit constraints can avoid unreasonable structures, they also shrink the search space and limit expressiveness
The right panel shows the EIC distribution under different noise levels. With EIC, both MCTS and GP consistently achieve lower values than their vanilla versions, approaching those of real physical formulas. This demonstrates that EIC not only improves performance but also prevents the generation of structurally unreasonable, hard-to-interpret formulas, making it a better constraint solution than unit constraints or manually designed nested rules.
@@
Figure 5: Pareto fronts and EIC distributions on the white-box dataset. $\dagger$ and $\ddagger$ indicate generative and regression methods, respectively, while others are heuristic methods. In the left panel, the lines show the Pareto front tiers, from bottom-left (best) to top-right (worst). In the right panel, markers indicate the mean EIC and its 95% confidence interval at different noise levels.
Black-box experimental results. Figure 6 shows the results on 122 black-box problems. In the left panel, incorporating EIC improved the Pareto front positions of GP and MCTS, moving them from Tier 2 and Tier 4 to Tier 1 and Tier 3, respectively, demonstrating that excluding structurally unreasonable formulas with EIC can enhance symbolic regression performance even on real-world problems without known ground truth. However, while EIC improved both accuracy and simplicity for MCTS, it improved simplicity but reduced accuracy for GP. This likely reflects the presence of datasets with relationships too complex to be captured by interpretable formulas of reasonable structure. As a regularizer, EIC guides search away from unreasonable forms, but also prevents generating overly complex formulas to overfit the intricate patterns in the data.
The right panel shows EIC distributions, where incorporating EIC substantially reduces unreasonable structures in both MCTS and GP. Compared to the white-box results in Figure 5, this effect is even stronger in the black-box setting, suggesting that when underlying relationships are unclear, symbolic regression tends to rely on overly complex functional forms, leaving more room for improvement in formula plausibility.
Figure 6: Pareto fronts and EIC distributions on the black-box dataset. * indicates decision-tree methods, other symbols are the same as in Fig 5.
4.3 Incorporate EIC into Generative Symbolic Regression Methods
Experimental setups. In this experiment, we considered E2ESR (kamienny2022end, ), SNIP (meidani2023snip, ), and SR4MDL (yu2025symbolic, ), covering the progression of this methodology from earlier efforts to more recent advances over the past two years. Each method was trained on randomly generated formulas and evaluated on the 133 SRBench white-box problems. For E2ESR and SNIP, we used the $R^{2}$ of generated formulas as the performance metric; for SR4MDL, which predicts minimum description length rather than formula tokens [^1] Strictly speaking, this is not a generative method. However, since it uses the same model architecture and data generation scheme as E2ESR and SNIP, we include it for comparison., we used the RMSE and MAE of its prediction as metrics. Training stopped when test performance plateaued over five million samples. We then retrain the models under the same hyperparameters using filtered formulas with $\text{EIC}<\theta$ until performance matches that obtained with unfiltered data. Based on Figure 2, we set $\theta=2$ to align the EIC of filtered formulas with that of real physical formulas (see Appendix A.3 for details).
@@ 4.3 Incorporate EIC into Generative Symbolic Regression Methods
We also evaluate a recent data construction baseline, PhyE2E (ying2025neural, ), which uses LLMs fine-tuned on physical equations to generate “look-physical” formulas with unit constraints for pretraining. We trained the E2ESR model on 180k PhyE2E formulas. To avoid unfair comparisons due to its limited sample size, we used a mixed sampling strategy that samples from PhyE2E formulas with a probability of 0.1 and generates random formulas with a probability of 0.9.
@@
EIC reduces the distribution gap between training samples and real physical formulas. We evaluated whether EIC filtering produces a training distribution closer to real-world formulas. We generated 1024 formulas from the unfiltered random generator and 1024 from the filtered generator, and compared them with three benchmark sets, including Feynman (119 physics formulas), Strogatz (14 ODE formulas), and Wiki Named Equations (984 equations, see Appendix B.2 for details). Similarity was measured via variable counts, constant counts, operator counts, and formula length, using Jensen-Shannon (JS) and Kullback-Leibler (KL) divergences.
@@
Figure 7: Test set performance of E2ESR trained on different samples. The grey line shows the convergence trained on random formulas.
EIC improves sample efficiency and generalization performance. The results are summarized in Table 2, where models trained on randomly generated formulas eventually generalized to the Feynman test set, but typically required tens of millions of samples. In contrast, training on EIC-filtered formulas achieved equal or better performance with fewer training samples, improving sample efficiency by 357%, 233%, and 287% on E2ESR, SNIP, and SR4MDL, respectively. This demonstrates that EIC filtering removes structurally unreasonable formulas, producing training data that is more physically meaningful and transferable to real-world tasks, thereby enhancing pretraining-based symbolic regression.
Figure 7 shows that training on EIC-filtered samples further improved final performance compared to random formulas. For E2ESR, training for the same number of steps increased $R^{2}$ from 0.5559 to 0.6808, representing a 22.4% relative gain. Similar improvements were observed for SNIP and SR4MDL, with relative gains of 13.5% and 5.14%, respectively (see Appendix B.2). Incorporating PhyE2E formulas into pretraining did not improve performance and slightly reduced generalization, likely due to a trade-off between unit consistency and formula diversity: although PhyE2E formulas satisfy unit constraints, their limited diversity reduces pretraining effectiveness. In contrast, EIC filtering preserves both structural plausibility and distributional diversity, thus yielding superior generalization.
Table 1: Distribution differences between real physical formulas and generated formulas. We generated 1024 random formulas and 1024 EIC-filtered formulas ($\mathrm{EIC}<2.0$) and compared them with three real physical formula sets.
@@
Table 2: Sample efficiency of different methods trained with random and filtered formulas.
Pre-training Formulas E2ESR SNIP SR4MDL
#Sample↓ R2↑ #Sample↓ R2↑ #Sample↓ MAE↓ RMSE↓
@@
Efficiency $\Delta(\%)$ 357.5% 198.8% 287.4%
4.4 EIC Reflects Interpretability in Alignment with Human Experts
Experimental setups. To test whether EIC aligns with human judgments of formula interpretability, we compared formulas discovered by EIC-MCTS and PySR, which were selected for their top performance in EIC and complexity, respectively (Section 4.2). We focused on Pareto fronts they discovered from 19 one- and two-dimensional PMLB black-box problems, where data can be visualized and understood by human experts. From each problem, we selected pairs of formulas with similar $R^{2}$ and complexity but differing EIC, yielding 172 pairs across all problems. We invite 108 volunteer participants with at least a bachelor’s degree in science or engineering, to whom we assign 10 pairs of formulas and ask to choose the more interpretable ones from each pair. We collected 840 valid evaluations, with each pair assessed on average 4.9 times (see Appendix B.3).
EIC evaluates formula interpretability in alignment with human experts. The results are summarized in Figure 14, where the left two panels show the distribution of complexity and $R^{2}$ for the selected formula pairs, indicating comparable performance between the two methods. The right two panels show that EIC-MCTS produces formulas with substantially lower EIC values than PySR, and human experts preferred EIC-MCTS in 70.12% of evaluations, which is 2.3 times more than PySR. This suggests EIC captures interpretability in a manner aligned with expert intuition. Moreover, when asking LLMs to act as domain experts, they similarly preferred EIC-MCTS 72.19% of the time, supporting the reliability of the rating results (see Appendix B.3).
Figure 8: Distributions of complexity, accuracy, EIC, and expert preference of formula pairs.

### v2

@@ 4 Experiments
In this section, we first reveal the structural stability gap in existing SR methods using EIC (Section 4.1). Then, we demonstrate the practical utility of EIC in enhancing search-based performance (Section 4.2), improving generative sample efficiency (Section 4.3), and validating its alignment with human expert preferences for interpretability (Section 4.4).
4.1 Structural Stability Gap in Existing SR Methods
@@
Experimental setups. To assess the structural rationality of existing SR discoveries, we benchmark 17 symbolic regression methods on 133 white-box problems from SRBench (119 Feynman and 14 Strogatz formulas), comparing the EIC distributions of the discovered formulas against the ground-truth physical formulas. The dataset of each problem is split into 75% training and 25% test data. Each method is executed 10 times per problem, with a filtering step that excludes poorly fitting results (Test set $R^{2}<0.8$) to ensure the analysis focuses on accurate candidates.
Figure 3: The structural stability gap between real physical formulas against SR discovered ones. The figure shows the average complexity and EIC of formulas discovered by 17 SR methods on 133 SRBench white-box problems (SBP-GP is not displayed due to its high EIC). The underfit formulas with test set $R^{2}$ below $0.8$ are ignored to focus on accurate candidates.
Figure 3 presents the average complexity and structural stability (EIC) of formulas discovered by 17 SR methods. Compared to the ground-truth physical formulas, the results reveal the following patterns regarding structural rationality:
Rigid constraints limit expressiveness. Methods such as DSR (Petersen et al., 2021) and RSRM (Xu et al., 2024) achieve the lowest EIC and complexity values, primarily due to their rigorous, manually designed constraints (e.g. forbidding nested trigonometric functions or mutually inverse operators). While these hard constraints effectively prune ill-conditioned structures, they overly restrict the search space, leading to inferior predictive accuracy (as we will demonstrate in the Pareto front analysis in Section 4.2). This highlights the necessity for a flexible, quantitative metric like EIC to guide the search rather than restricting it.
Difference between complexity and structural stability. Despite exhibiting the highest complexity, AIFeynman2 maintains a moderate EIC. We attribute this to physics-inspired strategies, such as divide-and-conquer and symmetry discovery, which build formulas from physically meaningful components. This confirms EIC identifies AI Feynman’s structural rationality, measuring intrinsic robustness rather than merely penalizing length.
Current objectives overlook structural stability. Apart from DSR and RSRM that overly restrict the search space, existing approaches like PySR (Cranmer, 2023) and SR4MDL (Yu et al., 2025), despite matching the compactness of the ground truth, fail to match its stability. Specifically, their EIC values ($\approx 2.0$) are significantly higher than real physical formulas (<1.0<1.0), indicating a difference in noise amplification of over an order of magnitude. This suggests that the standard bi-objective framework (balancing accuracy and complexity, or even absolute Pareto optimality (Fong and Motani, 2024)) is insufficient to prevent structurally irrational solutions, which motivates us to integrate EIC as a third objective to explicitly steer the search toward structurally rational regimes (as implemented in Section 4.2).
Generative priors bias toward structural instability. While generative methods like SNIP (Meidani et al., 2023) and NeurSR (Biggio et al., 2021) match search-based counterparts (PySR, SR4MDL) in complexity, they exhibit significantly higher EIC. This instability extends to Large Language Models (Appendix C.1), whose synthesized formulas (Shojaee et al., 2025) remain far less stable than physical ground truths. This suggests that pre-trained models, despite producing compact formulas, often hallucinate noise-amplifying substructures. We attribute this to the distributional shift between random pre-training data and physical reality, leading to poor generalization. Consequently, we employ EIC to filter pre-training corpora, aligning the training distribution with reality to boost sample efficiency and structural rationality (Section 4.3).
@@ 4.2 Incorporating EIC into Heuristic Search Methods
Figure 4: Pareto fronts on the SR benchmark. $\dagger$, $\ddagger$, *, and ** indicate generative, regression, decision-tree, and black-box methods, respectively, while others are search-based methods. The lines show the Pareto front tiers, from bottom-left (best) to top-right (worst).
Experimental setups. To evaluate EIC’s efficacy in guiding search, we integrated it into two representative heuristic algorithms: Genetic Programming (GP) (Augusto and Barbosa, 2000) and Monte Carlo Tree Search (MCTS) (Sun et al., 2022), which underpin widespread methods like PySR, TPSR (Shojaee et al., 2023), SR4MDL, etc. We modify the standard fitness function $f=\eta^{C}/(1+\text{NMSE})$, where $C$ denotes the formula complexity (length), $\text{NMSE}=\text{MSE}/\text{Var}(y)$, $\eta<1$ penalizes formula complexity (Sun et al., 2022), introducing an EIC penalty term: $f_{\text{new}}=f-\alpha\cdot\text{EIC}$, where $\alpha$ steers the search away from structurally irrational solutions (see Appendix B.1 for details). Notably, we employ linear regression to optimize constants during the search, avoiding more costly nonlinear optimization algorithms such as BFGS(Fletcher, 2000) and accelerating the search. We benchmark these EIC-enhanced variants against 17 baselines on SRBench, comprising 133 white-box and 122 black-box problems. For each problem, we conduct 10 independent trials with a 4-hour limit using a 75%/25% train-test split. For white-box tasks, we further assess robustness by injecting Gaussian noise at three levels $\{10^{-3},10^{-2},10^{-1}\}$. Performance is evaluated based on the Accuracy-Complexity Pareto frontier, which is the standard metric in SR. Besides, we also evaluate the results’ EIC to assess structural rationality.
White-box experimental results. Figure 4 (left) summarizes white-box results (full results in Appendix C.2). Integrating EIC advanced MCTS and GP from the second to the first Pareto tier, improving both their accuracy and compactness. This improvement can be attributed to EIC explicitly penalizing structurally irrational patterns, which prevents excessive operator stacking and acts as a regularizer to enhance generalization. Consequently, although EIC targets an objective distinct from the standard Accuracy-Complexity trade-off, it synergistically improves both metrics. This structural improvement is corroborated by Appendix Figure 9, where EIC values for both methods drop significantly toward ground-truth levels—a natural consequence of explicitly minimizing EIC during the search. In contrast, while DSR and RSRM yield compact formulas with low EIC (Figure 3), their significantly lower accuracy suggests that their manual constraints, while effectively avoiding unreasonable structures, inevitably shrink the search space and limit expressiveness.
Black-box experimental results. Figure 4 (right) presents results on 122 black-box problems (full results in Appendix C.2). Incorporating EIC advanced the Pareto fronts of GP (Tier 2 $\to$ 1) and MCTS (Tier 4 $\to$ 3), demonstrating that excluding structurally unreasonable formulas enhances performance even without ground truth. This is corroborated by significantly reduced EIC values in Appendix Figure 9. Notably, this reduction is more pronounced than in white-box settings, suggesting EIC effectively counters SR’s tendency to rely on overly complex forms when underlying relationships are opaque. However, while MCTS improved in both accuracy and simplicity, GP improved simplicity but slightly reduced accuracy. This is probably due to a trade-off between numerical approximation and structural rationality. Standard GP often exploits numerically unstable structures (e.g., high-degree polynomials) as universal approximators to minimize error, yielding accurate but physically meaningless formulas. EIC constrains the search to stable forms, purposely sacrificing the marginal accuracy gained from such unstable approximations to ensure physical plausibility.
4.3 Incorporate EIC into Generative SR Methods
Current generative approaches, such as E2ESR (Kamienny et al., 2022), SNIP (Meidani et al., 2023), and SR4MDL (Yu et al., 2025), typically rely on the random formula generation algorithm developed by Lample and Charton (2020) for pretraining. While this mechanism provides unlimited training data, the generated synthetic formulas often exhibit unstable structures rarely seen in real-world physical equations. Consequently, models pre-trained on such data suffer from severe sample inefficiency, often requiring tens of millions of samples to achieve effective generalization.
To address this, we propose an EIC-guided rejection sampling strategy to filter the pretraining corpus. Specifically, during data generation, we evaluate the EIC value of each sample and discard those exceeding a threshold $\theta$, regenerating them until the condition is met. This process removes structurally implausible samples while preserving the overall diversity of the training data. Based on the distribution of real physical formulas shown in Figure 2, we set $\theta=2$ to enforce structural alignment with physical reality (see Appendix B.2 for details).
Experimental setups. In this experiment, we applied this EIC-guided rejection sampling strategy to E2ESR, SNIP, and SR4MDL, covering the progression from early baselines to state-of-the-art methods. Each method was trained on both unfiltered (random) and filtered ($\text{EIC}<2$) formulas, and evaluated on the 133 SRBench white-box problems. Performance was measured using $R^{2}$ for E2ESR and SNIP, and RMSE/MAE for SR4MDL [^2] Strictly speaking, SR4MDL predicts minimum description length rather than tokens. However, as it shares the same architecture and data scheme as E2ESR, we include it for comparison.. Training on unfiltered data was conducted until test performance plateaued (approx. 5 million samples), serving as the baseline target to evaluate the sample efficiency gains of the EIC-filtered training.
@@ 4.3 Incorporate EIC into Generative SR Methods
Additionally, we benchmark against a recent data construction baseline, PhyE2E (Ying et al., 2025), which uses LLMs to generate “look-physical” formulas with unit constraints. We trained E2ESR on 180k PhyE2E formulas. To ensure a fair comparison, given its limited dataset size, we employed a mixed sampling strategy: sampling from PhyE2E formulas with a probability of 0.1 and generating random formulas with a probability of 0.9.
@@
EIC-guided filtering reduces the distribution gap between training samples and real physical formulas. We evaluated whether EIC filtering produces a training distribution closer to real-world formulas. We generated 1024 formulas from the unfiltered random generator and 1024 from the filtered generator, and compared them with three benchmark sets, including Feynman (119 physics formulas), Strogatz (14 ODE formulas), and Wiki Named Equations (984 equations, see Appendix C.3 for details). Similarity was measured via variable counts, constant counts, operator counts, and formula length, using Jensen-Shannon (JS) and Kullback-Leibler (KL) divergences.
@@
Figure 5: Generalization performance of E2ESR trained on different samples. The grey line shows the convergence trained on random formulas.
EIC-guided filtering boosts sample efficiency and generalization performance. The results are summarized in Table 2, where models trained on randomly generated formulas eventually converged on the Feynman test set, but typically required tens of millions of samples. In contrast, training on EIC-filtered formulas achieved equal or better performance with fewer training samples, improving sample efficiency by 357%, 233%, and 287% on E2ESR, SNIP, and SR4MDL, respectively. This demonstrates that EIC filtering removes structurally unreasonable formulas, producing training data that is more physically meaningful and transferable to real-world tasks, thereby enhancing pretraining-based symbolic regression.
Figure 5 shows that training on EIC-filtered samples further improved final performance compared to random formulas. For E2ESR, training for the same number of steps increased $R^{2}$ from $0.55$ to $0.68$, representing a 22.4% relative gain. Similar improvements were observed for SNIP and SR4MDL, with relative gains of 13.5% and 5.14%, respectively (see Appendix C.3). Incorporating PhyE2E formulas into pretraining did not improve performance and slightly reduced generalization, likely due to a trade-off between unit consistency and formula diversity: although PhyE2E formulas satisfy unit constraints, their limited diversity reduces pretraining effectiveness. In contrast, EIC filtering preserves both structural plausibility and distributional diversity, thus yielding superior generalization.
Table 1: Distribution differences between real physical formulas and generated ones. The JS and KL divergence between three real physical formula sets: Feynman ($n=119$), Strogatz ($n=14$), and Wiki Named Equations datasets ($n=984$), as well as 1024 random formulas and 1024 EIC-filtered ones ($\mathrm{EIC}<2.0$).
@@
Table 2: Sample efficiency of different methods trained with random and filtered formulas. #S is the number of samples.
Pre-training E2ESR SNIP SR4MDL
Formulas #S↓ $R^{2}$ #S↓ $R^{2}$ #S↓ MAE↓ RMSE↓
@@
$\Delta\text{Efficiency}$ 357.5% 198.8% 287.4%
4.4 EIC Aligns with Human Preferences for Interpretability
Experimental setups. To validate EIC’s alignment with human interpretability preferences, we conducted a controlled study comparing formula pairs with similar accuracy and complexity but distinct EIC values. We pooled formulas discovered by search-based symbolic regression methods across black-box problems (Section 4.2). To ensure intuitive assessment, we restricted analysis to 1D and 2D problems allowing clear visualization. From this pool, we rigorously selected pairs with indistinguishable fitting accuracy ($\Delta R^{2}\leq 0.02$) and complexity ($\Delta C\leq 2$), yet significant structural divergence ($|\Delta\text{EIC}|\geq 2$). We invited 108 science/engineering volunteers to evaluate 10 random pairs each, selecting the more interpretable formula based on visualized behavior and mathematical form. In total, 870 valid evaluations were collected (see Appendix C.4).
EIC Aligns with Human Expert Preferences for Interpretability. The results, summarized in Figure 6, demonstrate a strong concordance between EIC scores and expert preferences. Although the paired formulas possess comparable distributions of complexity and accuracy (left panels), human experts exhibited a decisive preference for the formulas with lower EIC scores, favoring them in 69.8% of evaluations (95% CI: [67%, 73%]). A one-sided binomial test confirms this preference is statistically significant far beyond random chance ($p<10^{-30}$). This result indicates that when accuracy and complexity are controlled, EIC serves as a critical third dimension that effectively captures the structural rationality preferred by humans. To assess the reliability of these subjective judgments, we calculated the inter-rater reliability using Fleiss’ Kappa, yielding $\kappa=0.37$. This indicates a fair level of agreement among experts, suggesting that EIC captures an objective structural quality recognized by humans rather than random individual preferences. Furthermore, Large Language Models (LLMs) acting as domain experts exhibited a similar preference pattern (favoring low-EIC formulas in 72.19% of cases), providing independent corroboration of the human rating results.
Figure 6: Comparison of formula pairs. “Lower” and “Higher” are formulas with the relatively lower/higher EIC values within each pair, whose complexity and accuracy ($R^{2}$) are indistinguishable but elicit a decisive expert preference for the lower candidates.